# si_topodynamics.pdf Dynamics of transport by helical edge states

### v1

@@ Abstract
Topologically nontrivial band structure of a material may give rise to special states that are confined to the material’s boundary and protected against disorder and scattering. Quantum spin Hall effect (QSHE) is a paradigmatic example of phenomenon in which such states appear in the presence of time-reversal symmetry in two dimensions. Whereas the spatial structure of these helical edge states has been largely studied, their dynamic properties are much less understood. We design a microwave experiment mimicking QSHE and explore the spatiotemporal dynamics of unidirectional transport of optical angular momentum (or pseudospin) by edge states. Pseudospin-polarized signal propagation is shown to be immune to scattering by defects introduced along the edge. Its velocity is two to three orders of magnitude slower than the speed of light in the free space, which may have important consequences for practical applications of topological edge states in modern optical and quantum-information technologies.
@@
Quantum Hall effect (QHE) [1, 2] as well as its variants—quantum spin Hall [3, 4] and quantum anomalous Hall [5, 6] effects (QSHE and QAHE)—are archetypal examples of topological phenomena in condensed matter physics. They are due to the nontrivial topological structure of electronic bands in some two-dimensional (2D) materials, leading to the appearance of robust conducting edge states protected against scattering by defects and giving rise to a quantized conductance [7, 8]. The direct relationship between edge states and quantized conductance has been challenged by recent experiments in which electric current density in QAHE has been measured with spatial resolution and found to be not restricted to sample edges [9, 10]. Together with theoretical modelling [11], such measurements unveil unexpected microscopic details of the robust conductance quantization in QAHE and demonstrate that our current understanding of topological phenomena in condensed matter physics is far from being complete.
@@
Whereas spatially resolved measurements represent a real experimental challenge in 2D electron gases, they are quite routinely performed in photonic or acoustic setups designed to mimic electronic systems. Indeed, classical-wave analogs of quantum Hall effects has been proposed and experimentally demonstrated rather soon after their discoveries, first in photonics [12, 13, 14, 15] and then in acoustics [16, 17]. Spatial maps of eigenmodes or wave fields in topologically nontrivial samples have been presented in virtually every publication in this research field. Cold-atom experiments in which the role of 2D electron gas is played by an ultracold atomic gas also allow for measurements with spatial resolution [18, 19, 20, 21]. In addition to facilitating position-resolved measurements, classical-wave and cold-atom systems have the advantage of being easier to control, allowing for observing the investigated physical phenomena in their purest form. In particular, interactions between particles (photons, phonons or atoms) can be made negligibly weak, disorder can be finely tuned, and finite-temperature effects almost fully suppressed. This makes such systems interesting to clarify details of transport phenomena in quantum Hall effects by going beyond the transverse and longitudinal conductivity measurements that are common in electron transport experiments.
@@
In contrast to the spatial structure of transport in topologically nontrivial systems, its dynamical aspects are largely unexplored to date. First measurements of propagation velocities associated with topologically protected edge states have been performed in ultracold quantum gases only very recently [20, 21]. These works deal with chiral edge states that arise in systems with broken time-reversal (TR) symmetry and can be roughly thought of as analogous to those in QHE or QAHE. Dynamical aspects of transport due to helical edge states in TR-symmetric QSHE-like systems have never been addressed up to now. In the present work, we fill this gap by performing experiments in a 2D microwave setup mimicking QSHE, see Fig. 1(a) and Supplemental Material (SM) 111 See SM that contains information on the experimental setup, details on the mapping between the experimental system and the tight-binding model, the computation of the group velocity in arrays of coupled oscillators, the computation of the angular momentum of light in six-cylinder clusters, and the computation of the topological invariants. The SM includes Refs. [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. Sec. 1. The role of spin—the intrinsic angular momentum of electrons—is played in our system by the orbital angular momentum of electromagnetic energy flow in hexagonal clusters (side ) of weakly coupled microwave resonators (dielecrtric cylinders) playing the role of elementary cells and arranged in a triangular lattice of spacing [24]. The symmetry behind topological phenomena in our experiments is the (pseudo-)TR one: an analog of TR operator is constructed as , where with the Pauli matrix, and is the complex conjugation [25]. Due to symmetry of the hexagonal six-cylinder structure, its - and -orbitals are eigenvectors of with , which produces Kramers degeneracy of bands at the analog of Dirac point in the band structure of our experimental system when and cylinders form a honeycomb lattice. Expanding six-cylinder clusters () opens a topological gap in the spectrum at Dirac point, with edge states crossing the gap in a sample with boundaries [25]. A narrow spectral gap remains open between the bands corresponding to edge states, due to the breakdown of symmetry near sample edges [24, 25], see Fig. 1(b). It barely affects our measurements because of the limited spectral resolution of the latter. We thus ignore this gap in the following. A number of previous experiments with microwaves [26, 27] and light [28, 29, 30] have already explored topological wave phenomena in similar structures. The novelty of our work resides in position- and time-resolved measurements of the pseudospin allowing us to directly observe unidirectional propagation of microwaves and pseudospin transport along an interface between topologically distinct parts of the sample. We demonstrate that the transport of pseudospin is not perturbed by defects introduced at the interface and measure both the group velocity of waves propagating along the interface and the velocity of pseudospin transport.
@@
In our experiments, microwave reflection and transmission via TE modes of the lattice of dielectric cylinders are measured using two loop antennas, one of which is fixed in the middle of the sample and the other one is placed at a position [31, 32, 33] (see Fig. 1(a) and SM [23] Sec. 1). The explored frequency range –7.7 GHz around the first magnetic-dipole resonance of an individual cylinder is below the cut-off of the first TE mode of the empty Fabry-Perot cavity hosting the sample (where is the speed of light in the free space), so that electromagnetic excitations only exist inside the cylinders and the latter are coupled through the spatial overlap of evanescent waves. Microwave propagation in such a system can be mapped on a tight-binding model with nearest-neighbor couplings that decreases exponentially with the distance between cylinders [34]. We define as the coupling between cylinders of the same elementary cell and between cylinders of different elementary cells (SM [23] Sec. 2). Wu and Hu have predicted the opening of a band gap in the spectrum of such a model when , corresponding to in our case [24, 25]. Figure 1(c) indeed shows that the density of states gets depressed around GHz when . The spectral range in which the depression takes place is well described by the tight-binding calculation yielding red lines in Fig. 1(c).
@@
Formally replacing the electron spin operator by a pseudospin corresponding to the orbital angular momentum of the electromagnetic wave in six-cylinder clusters allows for computing topological invariants by analogy with QSHE: the spin Chern number [35] and the spin Bott index [36] (see SM [23] Sec. 5). The results obtained for the tight-binding model on which our experimental system can be mapped are shown in Fig. 1(d). The topological character of the gap for predicted earlier [24, 25] is witnessed by (computed for the infinite lattice) and (computed for a lattice of finite size) that become different from zero. Note that the values of and computed using periodic boundary conditions (PBC) coincide exactly, which is in agreement with recent mathematical literature [37]. We also compute the spin Bott index for a lattice with open boundary conditions that correspond to the experimental situation. The result depends on the lattice size and becomes different from zero at which is slightly larger than , signaling that the finite sample size perturbs topological properties. This effect attenuates with increasing lattice size and convergence of towards is observed.
@@
The topological invariants shown in Fig. 1(d) are difficult to measure in an experiment. To observe an experimental manifestation of a topologically nontrivial band gap in our system, we create a sample consisting of two topologically different halves: , on the left and , on the right. The interface between the two halves of the sample should carry topologically protected edge states that we evidence by using the transmission coefficients to compute the response of the sample to a pulsed excitation (central frequency around 7.4 GHz) by a loop antenna placed at opposite ends of the interface (top or bottom of the interface for the first and second rows of Fig. 2, respectively). The spatiotemporal evolution of pseudospin (see SM [23] Sec. 4 for details) is shown in Supplemental videos 1–4 222See SM Videos showing the propagation of a Gaussian pulse between topologically distinct regions (videos 1–4) and around a defect (videos 5 and 6).. Here is the position of the cylinder with respect to the center of a six-cylinder cluster. Figure 2 shows the time integral of . We clearly see that in all cases, propagation takes place along boundaries between either topologically distinct parts of the sample (Figs. 2(a) and (c)) or the topologically nontrivial part and the free space (Figs. 2(b) and (d)). The sign of is locked to the direction of propagation: or for clockwise or counter-clockwise propagation, respectively. This is in full analogy with QSHE in which electrons with opposite spins propagate in opposite directions along an edge of a 2D sample. The direction in which a wave excited by a source at a given position propagates is controlled by the frequency of the source : for below (above) the middle of the gap, the system is on the decreasing (rising) part of the dispersion curve of the edge state and the propagation is counter-clockwise (clockwise) with (). The fact that a single edge mode dominates in all cases can be explained by an accurate analysis of the band diagram in Fig. 1(b), see SM [23] Sec. 2.3.
@@
In addition to yielding appealing visual representations of the dynamics of helical edge states (see Supplemental videos 1–4 [38]), our measurements allow for measuring their velocities. We define the group velocity as the velocity of propagation of the maximum of intensity . is shown by open green squares in Fig. 3 and agrees well with the theoretical prediction following from the simulation of the equivalent tight-binding model (black dashed line, see SM [23] Sec. 2.4). We also define the velocity of pseudospin transport , where is the barycenter of in a six-cylinder cluster centered at : , and is a unit vector parallel to the interface between topologically distinct parts of the sample. In principle, may be different from , which would signal that speeds of pseudospin and energy transports are not the same. However, our measurements yield within experimental accuracy, see Fig. 3. and exhibit very little dependence on , remaining roughly constant within the range — explored in the experiment. Furthermore, we observe that the measured velocities are of the same order of magnitude as the velocity predicted and observed in a 1D chain of coupled resonators, i.e. with mm and MHz (blue dash-dotted line in Fig. 3, see SM [23] Sec. 3). The same calculation done for coupled harmonic oscillators arranged in a honeycomb lattice yields velocities that are a factor of two larger (See SM [23] Sec. 3.2). This observation emphasizes the 1D character of the helical edge states.
@@
Robustness against scattering is a distinctive feature of edge states arising due to the topologically nontrivial band structure in the bulk [7, 8]. In order to demonstrate it in our system, we prepare a sample with a defect located along the boundary separating the two topologically distinct parts of the sample. The defect consists of two six-cylinder clusters of size protruding into the part of the sample where , see Fig. 4. Repeating the measurements reported in Fig. 2(a) and (c) produces results shown in Fig. 4 and Supplemental videos 5 and 6 [38]. As we see in Figs. 4(a) and (c), the defect does not perturb the unidirectional transport of pseudospin. Microwaves go around the defect without loosing their pseudospin polarization. Figures 4(b) and (d) demonstrate the absence of backscattering of microwaves by the defect: the temporal profiles of at the two six-cylinder clusters in front of the defect exhibit single maxima corresponding to the passage of the microwave pulse but do not show any secondary peak that might be caused by reflection of the wave by the defect.
@@
Topologically protected edge states in TR-symmetric photonic systems have been proposed as an efficient means to guide light in optoelectronic devices [39, 40, 41, 27]. The low velocity of propagation of pseudospin-polarized signal measured in our experiments imposes constraints on the information transfer rate of an optical communication channel using such states even though the precise value of can be adjusted according to the relation , which allows for tuning between and by varying the nearest-neighbor spacing in our experiment from 7 to 20 mm. Another possible application of topological edge states is their use as logical qubits that would be resilient to noise due to their nonlocal character [42, 43]. In this case, the finite speed of signal propagation limits the maximum rate at which such a qubit can be addressed because of nonzero time needed for the qubit to reach steady state. For a 2D logical qubit composed of elementary cells (a smaller system may not exhibit a proper band gap in the bulk), the time for a signal to go around the system is –5 s. This is already much longer than 25 ns needed to operate a single-qubit gate in the contemporary superconducting quantum processors [44]. The above examples show that the dynamical transport properties of topologically protected edge states may be crucial for design of their practical applications.
@@ Acknowledgements.
This work was funded by the Agence Nationale de la Recherche (Grant No. ANR-20-CE30-0003 LOLITOP) and supported by the French government through the France 2030 investment plan managed by the Agence Nationale de la Recherche as part of the Initiative of Excellence Université Côte d’Azur under reference number ANR-15-IDEX-01. The authors are grateful to the Université Côte d’Azur’s Center for High-Performance Computing (OPAL infrastructure) for providing resources and support.References
@@
- Yao et al. [2024] R. Yao, S. Chi, B. Mukherjee, A. Shaffer, M. Zwierlein, and R. J. Fletcher, Observation of chiral edge transport in a rapidly rotating quantum gas, Nat. Phys. (2024).

### v2

@@ Abstract
Topologically nontrivial band structure of a material may give rise to special states that are confined to the material’s boundary and protected against disorder and scattering. Quantum spin Hall effect (QSHE) is a paradigmatic example of phenomenon in which such states appear in the presence of time-reversal symmetry in two dimensions. Whereas the spatial structure of these helical edge states has been largely studied, their dynamic properties are much less understood. We design a microwave experiment mimicking QSHE and explore the spatiotemporal dynamics of unidirectional transport of optical angular momentum (or pseudospin) by edge states. Pseudospin-polarized signal propagation is shown to be immune to scattering by defects introduced along the edge. Its velocity is 2 to 3 orders of magnitude slower than the speed of light in the free space, which may have important consequences for practical applications of topological edge states in modern optical and quantum-information technologies.
@@
Quantum Hall effect (QHE) [1, 2] as well as its variants—quantum spin Hall [3, 4] and quantum anomalous Hall [5, 6] effects (QSHE and QAHE)—are archetypal examples of topological phenomena in condensed matter physics. They are due to the nontrivial topological structure of electronic bands in some two-dimensional (2D) materials, leading to the appearance of robust conducting edge states protected against scattering by defects and giving rise to a quantized conductance [7, 8]. The direct relationship between edge states and quantized conductance has been challenged by recent experiments in which electric current density in QAHE has been measured with spatial resolution and found to be not restricted to sample edges [9, 10]. Together with theoretical modeling [11], such measurements unveil unexpected microscopic details of the robust conductance quantization in QAHE and demonstrate that our current understanding of topological phenomena in condensed matter physics is far from being complete.
@@
Whereas spatially resolved measurements represent a real experimental challenge in 2D electron gases, they are quite routinely performed in photonic or acoustic setups designed to mimic electronic systems. Indeed, classical-wave analogs of quantum Hall effects have been proposed and experimentally demonstrated rather soon after their discoveries, first in photonics [12, 13, 14, 15] and then in acoustics [16, 17]. Spatial maps of eigenmodes or wave fields in topologically nontrivial samples have been presented in virtually every publication in this research field. Cold-atom experiments in which the role of 2D electron gas is played by an ultracold atomic gas also allow for measurements with spatial resolution [18, 19, 20, 21]. In addition to facilitating position-resolved measurements, classical-wave and cold-atom systems have the advantage of being easier to control, allowing for observing the investigated physical phenomena in their purest form. In particular, interactions between particles (photons, phonons or atoms) can be made negligibly weak, disorder can be finely tuned, and finite-temperature effects can be almost fully suppressed. This makes such systems interesting to clarify details of transport phenomena in quantum Hall effects by going beyond the transverse and longitudinal conductivity measurements that are common in electron transport experiments.
@@
In contrast to the spatial structure of transport in topologically nontrivial systems, its dynamical aspects are largely unexplored to date. First measurements of propagation velocities associated with topologically protected edge states have been performed in ultracold quantum gases only very recently [20, 21]. These works deal with chiral edge states that arise in systems with broken time-reversal (TR) symmetry and can be roughly thought of as analogous to those in QHE or QAHE. Studies of dynamical aspects of transport by helical edge states in TR-symmetric QSHE-like systems are scarce. A notable example of such a study demonstrated helical transport in a network of pairs of coupled mechanical oscillators [23]. In the present work, we perform experiments in a 2D microwave setup mimicking QSHE, see Fig. 1 (a) and Supplemental Material (SM) 111See SM that contains information on the experimental setup, details on the mapping between the experimental system and the tight-binding model, the computation of the group velocity in arrays of coupled oscillators, the computation of the angular momentum of light in six-cylinder clusters, and the computation of the topological invariants. The SM includes Refs. [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Sec. I. In contrast to Ref. [23] where an intricate and purposely designed coupling between oscillators was required, our setup relies on natural electromagnetic coupling between dielectric resonators. The role of spin—the intrinsic angular momentum of electrons—is played by the orbital angular momentum of electromagnetic waves. It originates from the microwave energy flow in hexagonal clusters (side ) of weakly coupled microwave resonators (dielectric cylinders) playing the role of elementary cells and arranged in a triangular lattice of spacing [25]. The symmetry behind topological phenomena in our experiments is the (pseudo-)TR one: an analog of the TR operator is constructed as , where with being the Pauli matrix, and is the complex conjugation [26]. Due to the symmetry of the hexagonal six-cylinder structure, its and orbitals are eigenvectors of with , which produces Kramers degeneracy of bands at the analog of the Dirac point in the band structure of our experimental system when and cylinders form a honeycomb lattice. Expanding six-cylinder clusters () opens a topological gap in the spectrum at the Dirac point, with edge states crossing the gap in a sample with boundaries [26]. Breakdown of symmetry near sample edges spoils the analogy between the pseudospin and the electron spin, leading to the opening of a narrow spectral mini-gap between the bands corresponding to edge states [see Fig. 1 (b)] and to the mixing of the two pseudospin channels [25, 26]. As convincingly demonstrated in previous experiments with microwaves [27, 28] and light [29, 30, 31], this does not compromise the topological nature of the band structure or the helicity of edge modes but limits the distance over which pseudospin-polarized propagation along sample edges can be observed. The focus of our work resides in position- and time-resolved measurements of the pseudospin , allowing us to directly observe unidirectional propagation of microwaves and pseudospin transport along an interface between topologically distinct parts of the sample. We demonstrate that the transport of pseudospin is not perturbed by defects introduced at the interface and measure both the group velocity of waves propagating along the interface and the velocity of pseudospin transport.
@@
In our experiments, microwave reflection and transmission via TE modes of the lattice of dielectric cylinders are measured using two loop antennas, one of which is fixed in the middle of the sample and the other one is placed at a position [32, 33, 34] (see Fig. 1 (a) and SM [24] Sec. I). The explored frequency range –7.7 GHz around the first magnetic-dipole resonance of an individual cylinder is below the cutoff of the first TE mode of the empty Fabry-Perot cavity hosting the sample (where is the speed of light in the free space), so that electromagnetic excitations only exist inside the cylinders and the latter are coupled through the spatial overlap of evanescent waves. Microwave propagation in such a system can be mapped on a tight-binding model with nearest-neighbor couplings that decrease exponentially with the distance between cylinders [35]. We define as the coupling between cylinders of the same elementary cell and between cylinders of different elementary cells (see SM [24] Sec. II). Wu and Hu have predicted the opening of a band gap in the spectrum of such a model when , corresponding to in our case [25, 26]. Figure 1 (c) indeed shows that the density of states gets depressed around GHz when . The spectral range in which the depression takes place is well described by the tight-binding calculation yielding red lines in Fig. 1 (c).
@@
Formally replacing the electron spin operator by a pseudospin corresponding to the orbital angular momentum of the electromagnetic wave in six-cylinder clusters allows for computing topological invariants by analogy with QSHE: the spin Chern number [36] and the spin Bott index [37] (see SM [24] Sec. V). The results obtained for the tight-binding model on which our experimental system can be mapped are shown in Fig. 1 (d). The topological character of the gap for predicted earlier [25, 26] is witnessed by (computed for the infinite lattice) and (computed for a lattice of finite size) that become different from zero. Note that the values of and computed using periodic boundary conditions (PBC) coincide exactly, which is in agreement with recent mathematical literature [38]. We also compute the spin Bott index for a lattice with open boundary conditions (OBC) that correspond to the experimental situation. The result depends on the lattice size and becomes different from zero at which is slightly larger than , signaling that the finite sample size perturbs topological properties. This effect attenuates with increasing lattice size and convergence of towards is observed.
@@
The topological invariants shown in Fig. 1 (d) are difficult to measure in an experiment. To observe an experimental manifestation of a topologically nontrivial band gap in our system, we create a sample consisting of two topologically different halves: , on the left and , on the right. The interface between the two halves of the sample should carry topologically protected edge states that we evidence by using the transmission coefficients to compute the response of the sample to a pulsed excitation (central frequency around 7.4 GHz) by a loop antenna placed at opposite ends of the interface (top or bottom of the interface for the first and second rows of Fig. 2, respectively). The spatiotemporal evolution of pseudospin (see SM [24] Sec. IV for details) is shown in Supplemental videos 1–4 222See SM Videos showing the propagation of a Gaussian pulse between topologically distinct regions (videos 1–4) and around a defect (videos 5 and 6).. Here is the position of the cylinder with respect to the center of a six-cylinder cluster. Figure 2 shows the time integral of . We clearly see that in all cases, propagation takes place along boundaries between either topologically distinct parts of the sample (Figs. 2 (a) and (c)) or the topologically nontrivial part and the free space (Figs. 2 (b) and (d)). The sign of is locked to the direction of propagation: or for clockwise or counter-clockwise propagation, respectively. This is in full analogy with QSHE in which electrons with opposite spins propagate in opposite directions along an edge of a 2D sample. The direction in which a wave excited by a source at a given position propagates is controlled by the frequency of the source : for below (above) the middle of the gap, the system is on the decreasing (rising) part of the dispersion curve of the edge state and the propagation is counter-clockwise (clockwise) with (). This is in contrast with the proposal of Ref. [25] where the use of a pseudospin-polarized source has been suggested to break the symmetry between modes with positive and negative whereas our source is equally efficient in exciting both. The fact that a single edge mode dominates in all cases can be explained by an accurate analysis of the band diagram in Fig. 1 (b), see SM [24] Sec. II.C.
@@
In addition to yielding appealing visual representations of the dynamics of helical edge states (see Supplemental Videos 1–4 [39]), our measurements allow for measuring their velocities. We define the group velocity as the velocity of propagation of the maximum of intensity . is shown by open green squares in Fig. 3 and agrees well with the theoretical prediction following from the simulation of the equivalent tight-binding model (black dashed line, see SM [24] Sec. II.D). We also define the velocity of pseudospin transport , where is the barycenter of in a six-cylinder cluster centered at : , and is a unit vector parallel to the interface between topologically distinct parts of the sample. In principle, may be different from , which would signal that speeds of pseudospin and energy transports are not the same. However, our measurements yield within experimental accuracy, see Fig. 3. and exhibit very little dependence on , remaining roughly constant within the range — explored in the experiment. Furthermore, we observe that the measured velocities are of the same order of magnitude as the velocity predicted and observed in a 1D chain of coupled resonators, i.e. with mm and MHz (blue dash-dotted line in Fig. 3, see SM [24] Sec. III). The same calculation done for coupled harmonic oscillators arranged in a honeycomb lattice yields velocities that are a factor of 2 larger (see SM [24] Sec. III.B). This observation emphasizes the 1D character of the helical edge states.
@@
Robustness against scattering is a distinctive feature of edge states arising due to the topologically nontrivial band structure in the bulk [7, 8]. In order to demonstrate it in our system, we prepare a sample with a defect located along the boundary separating the two topologically distinct parts of the sample. The defect consists of two six-cylinder clusters of size protruding into the part of the sample where , see Fig. 4. Repeating the measurements reported in Fig. 2 (a) and (c) produces results shown in Fig. 4 and Supplemental Videos 5 and 6 [39]. As we see in Figs. 4 (a) and (c), the defect does not perturb the unidirectional transport of pseudospin. Microwaves go around the defect without loosing their pseudospin polarization. Figures 4 (b) and (d) demonstrate the absence of backscattering of microwaves by the defect: the temporal profiles of at the two six-cylinder clusters in front of the defect exhibit single maxima corresponding to the passage of the microwave pulse but do not show any secondary peak that might be caused by reflection of the wave by the defect (see also SM [24] Sec. VI).
@@
Topologically protected edge states in TR-symmetric photonic systems have been proposed as an efficient means to guide light in optoelectronic devices [40, 41, 42, 28]. The low velocity of propagation of pseudospin-polarized signal measured in our experiments imposes constraints on the information transfer rate of an optical communication channel using such states even though the precise value of can be adjusted according to the relation , which allows for tuning between and by varying the nearest-neighbor spacing in our experiment from 7 to 20 mm. Another possible application of topological edge states is their use as logical qubits that would be resilient to noise due to their nonlocal character [43, 44]. In this case, the finite speed of signal propagation limits the maximum rate at which such a qubit can be addressed because of nonzero time needed for the qubit to reach steady state. For a 2D logical qubit composed of elementary cells (a smaller system may not exhibit a proper band gap in the bulk), the time for a signal to go around the system is –5 µs. This is already much longer than the 25 ns needed to operate a single-qubit gate in the contemporary superconducting quantum processors [45]. The above examples show that the dynamical transport properties of topologically protected edge states may be crucial for the design of their practical applications.
@@ Acknowledgements.
This work was funded by the Agence Nationale de la Recherche (Grant No. ANR-20-CE30-0003 LOLITOP) and supported by the French government through the France 2030 investment plan managed by the Agence Nationale de la Recherche as part of the Initiative of Excellence Université Côte d’Azur under Reference No. ANR-15-IDEX-01. The authors are grateful to the Université Côte d’Azur’s Center for High-Performance Computing (OPAL infrastructure) for providing resources and support.References
@@
- Yao et al. [2024] R. Yao, S. Chi, B. Mukherjee, A. Shaffer, M. Zwierlein, and R. J. Fletcher, Observation of chiral edge transport in a rapidly rotating quantum gas, Nat. Phys. 20, 1726 (2024).
@@
- Süsstrunk and Huber [2015] R. Süsstrunk and S. D. Huber, Observation of phononic helical edge states in a mechanical topological insulator, Science 349, 47 (2015).